A reaction is observed to have a rate constant of 35 min-1. Calculate the corresponding half-life for the reaction.


1
Expert's answer
2021-10-28T03:02:21-0400

Half life of first order reaction is


t1/2 = 0.693/K where K is rate constant = 25/min (Unit of rate constant for first order is S-1 and here given in Min-1 so it will be converted into S-1)

25 /min-1 = 25/60 S-1 = 0.42 S-1

Now putting the value K into the formula

t1/2 = 0.693/0.42 = 1.65

t1/2 = 1.65